OpenStudy (anonymous):

Evaluate 48 + 24 ÷ 8 – 2^2 47 51 55 64

OpenStudy (uri):

BODMAS or PEDMAS.

OpenStudy (anonymous):

Simplify the following: 48+24/8-2^2 Reduce 24/8 to lowest terms. The gcd of 24 and 8 is 8, so 24/8 = (8×3)/(8×1) = 8/8×3 = 3: 48+3-2^2 Evaluate 2^2. 2^2 = 4: 48+3-4 Evaluate 48+3 using long addition. | 1 | | 4 | 8 + | | 3 | 5 | 1: 51-4 Subtract 4 from 51. | 4 | 11 | 5 | 1 - | | 4 | 4 | 7: